# Fighting Exponentially Small Gaps by Counterdiabatic Driving

### v2

@@ Abstract
We investigate the efficiency of approximate counterdiabatic driving (CD) in accelerating adiabatic passage through a first-order quantum phase transition. First, we analyze a minimal spin-glass bottleneck model that is analytically tractable and exhibits both an exponentially small gap at the transition point and a change in the ground state that involves a macroscopic rearrangement of spins. Using the variational Floquet-Krylov expansion to construct CD terms, we find that while the formation of excitations is significantly suppressed, achieving a fully adiabatic evolution remains challenging. Extending our investigation to realistic NP-hard spin-glass problems—specifically, the $3$-regular Max Cut and $3$-XORSAT—we find again that local CD expansions lead to negligible improvements in the final ground state fidelity. These results highlight the limited impact of local CD methods in overcoming the bottlenecks associated with first-order quantum phase transitions. To address this limitation, we propose an alternative method, termed quantum brachistochrone counterdiabatic driving (QBCD), which employs the approximate full CD connecting the ground state and the first excited state at a single parameter value close to the critical point. In the minimal spin-glass model, QBCD enables exponentially faster adiabatic evolution than the local strategies. To bring experimental realization within reach for realistic NP-hard problems, we further enhance the QBCD Hamiltonian by sparsifying it to the density of the quadratic local expansion. This reduces the non-locality exponentially while preserving finite ground-state fidelity at driving times exponentially shorter than in the local strategies and in the counterdiabatic optimized local driving (COLD). These findings highlight that targeting a single avoided crossing with an exponentially reduced counterdiabatic correction based on approximate numerical knowledge is sufficient to bypass spin-glass bottlenecks, offering a compelling path forward for near-term quantum devices.
@@ I.1 Summary of results
In Sec. V, we introduce a novel approach to the bottleneck problem by attacking directly the first-order phase transition. The QBCD Hamiltonian is constructed using approximate knowledge of the eigenstates at a single parameter value at the transition. In the minimal spin glass model, the QCDB can be understood as driving the localized edge states associated with the minimal gap. We show that this approach achieves exponentially shorter adiabatic time scales by exploiting only an exponentially small fraction of the full counterdiabatic Hamiltonian. Combined with its system-size-independent energetic cost, these features support near-term experimental implementation.
We then test the generality of our findings in two realistic spin glass models representing NP-hard optimization problems, the Max Cut and the $3$-XORSAT. First, we demonstrate the same limited efficiency of the first and second-order local expansion via the ground state gap and the final fidelity. Next, we introduce a yet further enhanced version of the QBCD Hamiltonian, by sparsifying it to the density of the first order local expansion. Remarkably, this approach with an exponentially reduced non-locality provides the same level of shortening of the adiabatic time-scale as the full one. These features significantly ease the challenges of implementation, making the approach promising for both experimental platforms and large-scale classical simulations. Finally, we also test the efficiency of the COLD approach and find that while it offers a noticeable improvement over local expansion at small system sizes, this advantage gradually decreases as $L$ increases. We show that the sparsified QBCD clearly outperforms the COLD protocol. In addition, the numerical implementation of the COLD protocol is orders of magnitude more demanding than that of the other approaches, restricting its applicability to very small system sizes ($L\lesssim 10$).
@@ II Minimal model of spin-glass bottleneck
From the pictorial representation in Fig. 1a, it can be seen that Eq. (1) describes an Ising chain with uniform ferromagnetic couplings, except for three weaker couplings placed in diametrically opposite positions: two weak ferromagnetic ones ($J$) next to each other, and one weaker antiferromagnetic one ($-J^{\prime}$) on the other side. The condition $J^{2}<J^{\prime}$ in Eq. (3) is more technical, and is explained in App. B.4.
The Hamiltonian in Eq. (1) has $\mathbb{Z}_{2}$ fermion parity symmetry (the fermionic nature of the parity will become clear in Sec. II.2), due to the possibility of flipping all the spins along $z$:
$\Pi_{\mathrm{F}}=\prod_{j=1}^{L}\sigma_{j}^{x},\qquad\big{[}\Pi_{\mathrm{F}},H[\lambda]\big{]}=0.$ (4)
In the following, it will be convenient to restrict to the even parity sector, which is the one dynamically accessible starting from the ground state at $\lambda=0$.
The model has also $\mathbb{Z}_{2}$ reflection parity symmetry $\Pi_{\mathrm{R}}$, defined by the action
$\Pi_{\mathrm{R}}\sigma_{j}^{\alpha}\Pi_{\mathrm{R}}=\sigma_{L-j+1}^{\alpha},\qquad\alpha=x,y,z,$ (5)
and highlighted with a grey dashed line in Figs. 1a and 2b. One can check that $\big{[}\Pi_{\mathrm{R}},H[\lambda]\big{]}=0$. It will prove convenient to use a modified version of this symmetry to exploit the free fermionic nature of the chain (see Sec. II.2).
@@ II.1 Simple limits
In order to gain an understanding of the model, we analyze the lowest energy states at the beginning ($\lambda=0$) and at the end ($\lambda=1$) of the driving schedule. When $\lambda=0$, the ground state is simply the product state of spins pointing along the $x$ direction, |$\ket{E_{0}[\lambda=0]}=\prod_{j}(\ket{\uparrow}_{j}+\ket{\downarrow}_{j})/\sqrt{2}$. It has an even fermionic parity: $\Pi_{\mathrm{F}}\ket{E_{0}[\lambda=0]}=\ket{E_{0}[\lambda=0]}$, since the state on each site is even itself. The $L$ degenerate first-excited states are obtained by flipping one of the spins.
At the end of the driving schedule ($\lambda=1$), because of the presence of a single antiferromagnetic bond in a closed loop geometry, the ground state is frustrated. The condition $10<J^{\prime}<J<$ 1, Eq. (3), implies that the frustrated bond is at $j=2\ell+1$, see Fig. 1b, and thus the ground-state energy is $E_{0}=-(L-3)-2J+J^{\prime}$. Moreover, because of the fermion parity $\Pi_{\mathrm{F}}$, the ground state is doubly degenerate: the degeneracy is removed by restricting to the even fermion parity sector and taking the symmetric combination |$\ket{E_{0}^{+}}$ as in Fig. 1e.
Also the four degenerate first-excited states are easy to determine: they display a single frustrated bond, i.e.,  the weak ferromagnetic link $j=\ell$ or $\ell+1$, and have energies $E_{1}=-(L-3)-J^{\prime}$; see Fig. 1c–d for a pictorial representation (the two other degenerate states are obtained by flipping all the spins). The restriction to the even parity sector implies that symmetric combinations need to be taken, similarly to what is shown in Fig. 1e for the ground state.
From the exact computation reported below, it turns out that at an intermediate value of $\lambda$ the ground state undergoes an exponentially small avoided crossing with the first excited state. In particular, the ground state changes from a dressed version of the states in Fig. 1c–d ($\lambda<\lambda_{c}$) to a dressed version of the state in Fig. 1b ($\lambda>\lambda_{c}$). The ground states before and after the avoided crossing are macroscopically different, as the flipping of half of the chain is involved. Therefore, the Hamiltonian  (1) features a first-order transition similar to the ones that take place in spin-glass models.
The rest of this section is devoted to the analytical description of the avoided crossing of the model. While the computation was already performed in the original work [99], some steps were either missing or wrong (while the final results are substantially correct). For this reason, and to make the manuscript self-contained, we detail the solution and the salient features of the model.
@@ II.2 Fermionic representation and reflection parity symmetry
By means of a Jordan-Wigner transformation, the spin chain Eq. (1) can be fermionized into a chain of Dirac fermions, as detailed in App. A.1. The fermion parity operator $\Pi_{\mathrm{F}}$ in Eq. (4) is transformed into
$\Pi_{\mathrm{F}}=\prod_{j=1}^{L}\left(1-2c_{j}^{\dagger}c_{j}^{\phantom{\dagger}}\right),$ (6)
which counts the parity of the total number of fermions. Then, the Hamiltonian is transformed into
$H[\lambda]=\sum_{j=1}^{L}\left[-\lambda\tilde{J}_{j}\left(c_{j}^{\dagger}-c_{j}^{\phantom{\dagger}}\right)\left(c_{j+1}^{\dagger}+c_{j+1}^{\phantom{\dagger}}\right)\right.\\ \left.+2(1-\lambda)c_{j}^{\dagger}c_{j}^{\phantom{\dagger}}\right]-(1-\lambda)L,$ (7)
where the couplings $\tilde{J}_{j}$ are the same as $J_{j}$, except that the sign of the last coupling is flipped: $\tilde{J}_{L}=-J_{L}$. This is to take into account the even fermionic parity of the ground state sector; see App. A.1 for a thorough discussion.
Because of the particle non-conservation entailed by the Dirac Hamiltonian (7), it is convenient to pass to the Bogoliubov basis, where excitations are conserved in the adiabatic limit. However, the matrix representation of Eq. (7) is not analytically diagonalizable to the best of our knowledge (see also App. E.1). It is thus convenient to perform two additional transformations [99]. First, each Dirac fermion is split into a couple of Majorana fermions, as illustrated in Fig. 2a, according to:
$\gamma_{2j-1}:=\frac{1}{\sqrt{2}}\left(c_{j}^{\phantom{\dagger}}+c_{j}^{\dagger}\right),\quad\gamma_{2j}:=\frac{i}{\sqrt{2}}\left(c_{j}^{\phantom{\dagger}}-c_{j}^{\dagger}\right),$ (8)
that obey commutation relations $\{\gamma_{k},\gamma_{k^{\prime}}\}=\delta_{k,k^{\prime}}$. The Hamiltonian acquires the form
$H[\lambda]=-2i\sum_{j=1}^{L}\big{[}\lambda\tilde{J}_{j}\gamma_{2j}\gamma_{2j+1}+(1-\lambda)\gamma_{2j-1}\gamma_{2j}\big{]},$ (9)
see also Fig. 2a and App. A.2. Second, it is convenient to define a modified reflection symmetry $\widetilde{\Pi}_{\mathrm{R}}$, instead of $\Pi_{\mathrm{R}}$ in Eq. (5) (for a detailed explanation see App. A.3). Defining $\widetilde{\Pi}_{\mathrm{R}}$ by the action
$\widetilde{\Pi}_{\mathrm{R}}\gamma_{k}\widetilde{\Pi}_{\mathrm{R}}=\gamma_{2L-k+1},$ (10)
one can check that $\widetilde{\Pi}_{\mathrm{R}}^{2}=\mathbb{I}$ and
$\widetilde{\Pi}_{\mathrm{R}}H[\lambda]\widetilde{\Pi}_{\mathrm{R}}=-H[\lambda],$ (11)
i.e., that $\widetilde{\Pi}_{\mathrm{R}}$ is a chiral symmetry for $H[\lambda]$. This means that $H[\lambda]$ is not split into sectors and does not acquire a block diagonal form, but simply the spectrum of $H[\lambda]$ is symmetric with respect to 00, with eigenstates coming in pairs [100].
The presence of the chiral symmetry $\widetilde{\Pi}_{\mathrm{R}}$ suggests to define the symmetrized operators illustrated in Fig. 2b,
$\Gamma^{\pm}_{k}$ =±$=\frac{\pm i\gamma_{k}+\gamma_{2L-k+1}}{\sqrt{2}},\quad$ $k=1,3,\dots,L,$ (12a)
$\Gamma^{\pm}_{k}$ $=\frac{\gamma_{k}\pm i\gamma_{2L-k+1}}{\sqrt{2}},\quad$ $k=2,4,\dots,L-1,$ (12b)
where the factors ±$\pm i$ where chosen so that
$\{\Gamma_{k}^{+},\Gamma_{k^{\prime}}^{-}\}=\delta_{k,k^{\prime}},\quad\{\Gamma_{k}^{+},\Gamma_{k^{\prime}}^{+}\}=\{\Gamma_{k}^{-},\Gamma_{k^{\prime}}^{-}\}=0,$ (13a)
$\widetilde{\Pi}_{\mathrm{R}}\Gamma_{k}^{\pm}\widetilde{\Pi}_{\mathrm{R}}=\pm i\Gamma_{k}^{\mp},\quad(\Gamma^{\pm}_{k})^{\dagger}=\Gamma^{\mp}_{k}.$ (13b)
The equations above show that $\Gamma_{k}^{\pm}$ represent a new set of Dirac fermions, in terms of which the Hamiltonian takes the convenient form
@@
as shown in App. A.4. Above, the $J_{j}$’s have been restored since the last coupling is isolated from the others, and there is no need to keep the $\tilde{J}_{j}$’s. One can check again that $\widetilde{\Pi}_{\mathrm{R}}H[\lambda]\widetilde{\Pi}_{\mathrm{R}}=-H[\lambda]$, as follows immediately from Eqs. (13a)–(13b). Introducing the vectors
@@
From the explicit matrix above, one can see that the diagonalization of $H[\lambda]$ has been reduced to the diagonalization of a simpler one-particle hopping model with open boundary conditions, described by the matrix $\mathcal{H}[\lambda]$. Here and throughout the rest of the paper, we use calligraphic letters to refer to the operators acting on the linear space of the effective Hamiltonian $\mathcal{H}[\lambda]$, and a modified ket notation $|\psi)$ for the corresponding single-particle modes.
@@ II.3 Description of the avoided crossing
The matrix $\mathcal{H}[\lambda]$ in Eq. (16) cannot be diagonalized analytically for general $J_{j}$’s. However, using the $J_{j}$’s given in Eq. (2), the entire spectrum can be obtained with a Fourier ansatz: the effective hopping model is translation invariant in the bulk. Thus, the eigenstates must have a plane (or evanescent) wave structure in the bulk, and the effect of the inhomogeneous boundaries can be taken into account by simply fixing the boundary conditions. This is reminiscent of the solution of the Schrödinger equation on a line with a piecewise constant (or delta-function) potential. The computation is carried on in detail in App. B, while here, only the main features are stated.
Figure 3: (a) Sketch of the boundary-localized eigenmodes $|\psi_{\mathrm{L}})$,$|\psi_{\mathrm{R}})$ of the effective hopping Hamiltonian (16). (b) Spectrum of the hopping Hamiltonian (16) for $\ell=8$, $J=0.5$, $J^{\prime}=0.27$. The many-body ground state of the $\Gamma$ Hamiltonian in Eq. (14) is formed by populating all the negative-energy modes at $\lambda=0$ (solid lines) and leaving the positive-energy modes empty (dashed lines). The two localized modes $\epsilon_{\mathrm{L}}$ (blue) and $\epsilon_{\mathrm{R}}$ (orange) undergo an avoided crossing at $\lambda=\lambda_{\mathrm{c}}$: while $\epsilon_{\mathrm{R}}$ contributes to the ground state for $\lambda<\lambda_{\mathrm{c}}$ (continuous portion of the orange line), it is replaced by $\epsilon_{\mathrm{L}}$ for $\lambda>\lambda_{\mathrm{c}}$ (the continuous portion of the blue line). Since the spatial overlap between $|\psi_{\mathrm{L}})$ and $|\psi_{\mathrm{R}})$ is exponentially small in $L$, also the gap $\Delta_{\mathrm{min}}$ remains exponentially small according to Eq. (19).
The spectrum of the effective hopping model, Eq. (16), is essentially composed of plane wave eigenmodes. Under the condition $J^{2}<J^{\prime}$. However, a couple of boundary modes are present; see also Eq. (107): one, call it $|\psi_{\mathrm{L}})$, is localized around the left end $j=0$ of the $\Gamma$ chain, and a second one, denoted by $|\psi_{\mathrm{R}})$, extends around the right end $j=L$. Their spatial decay is regulated by the rates $\kappa_{\mathrm{L}}$ and $\kappa_{\mathrm{R}}$ as
@@
see Fig. 3a for a sketch. The corresponding eigenvalues $\epsilon_{\mathrm{L}}$ and $\epsilon_{\mathrm{R}}$ undergo an avoided crossing at a value of the driving parameter $\lambda=\lambda_{\mathrm{c}}$, with a gap
$\Delta_{\mathrm{min}}\propto e^{-\alpha L},\qquad\alpha=\frac{1}{2}\ln\left[\frac{J^{\prime}(1-J^{2})}{J^{2}-(J^{\prime})^{2}}\right].$ (19)
Notice that $\alpha>0$, thanks to the conditions Eq. (3). The exponentially small gap $\Delta_{\mathrm{min}}$ hinders adiabatic driving because the many-body ground state of the spin chain contains the single-particle orbital $|\psi_{\mathrm{R}})$ for $\lambda\lesssim\lambda_{\mathrm{c}}$, while it contains the single-particle orbital $|\psi_{\mathrm{L}})$ for $\lambda\gtrsim\lambda_{\mathrm{c}}$, see Fig. 3b. Therefore, the minimal gap along the adiabatic path is equal to $\Delta_{\mathrm{min}}$. At the avoided crossing, it is necessary to bring an effective particle from one end of the $\Gamma$ chain to the other end: in the language of spins, this corresponds to flipping an extensive number of them.

### v3

@@ Abstract
We investigate the efficiency of approximate counterdiabatic driving (CD) in accelerating adiabatic passage through exponentially small gaps. First, we analyze a minimal spin-glass bottleneck model that is analytically tractable and exhibits both an exponentially small gap at the transition point and a change in the ground state that involves a macroscopic rearrangement of spins. Using the variational Floquet-Krylov expansion to construct CD terms, we find that while the formation of excitations is significantly suppressed, achieving a fully adiabatic evolution remains challenging. Extending our investigation to realistic NP-hard spin-glass problems—specifically, the $3$-regular Max Cut and $3$-XORSAT—we find again that local CD expansions lead to negligible improvements in the final ground state fidelity. These results highlight the limited impact of local CD methods in overcoming the bottlenecks associated with first-order quantum phase transitions. To address this limitation, we propose an alternative method, termed quantum brachistochrone counterdiabatic driving (QBCD), which employs the approximate full CD connecting the ground state and the first excited state at a single parameter value close to the critical point. In the minimal spin-glass model, QBCD enables exponentially faster adiabatic evolution than the local strategies. To alleviate the challenges of its experimental and classical implementation for realistic NP-hard problems, we exponentially reduce the non-locality of the QBCD Hamiltonian by sparsifying its matrix elements to the density of the local expansions. Despite this drastic simplification, sparsified QBCD maintains finite ground-state fidelity at driving times exponentially shorter than in local strategies and counterdiabatic optimized local driving (COLD), demonstrating that even an exponentially reduced fraction of non-locality near the critical point is sufficient to achieve significant speedup across spin-glass bottlenecks. At the same time, reduced non-locality facilitates classical implementations via advanced numerical methods, while on quantum hardware its truncation serves as a resource-efficient approximation scheme for higher-order CD realizations.
@@ I.1 Summary of results
In Sec. V, we introduce a novel approach to the bottleneck problem by addressing directly the closing of the gap. The QBCD Hamiltonian is constructed using approximate knowledge of the eigenstates at a single parameter value near the transition. In the minimal spin glass model, the QCDB can be understood as driving the localized edge states associated with the minimal gap. We show that this approach achieves exponentially shorter adiabatic time scales by exploiting only an exponentially small fraction of the full counterdiabatic Hamiltonian. Combined with its system-size-independent energetic cost, these features support near-term experimental implementation.
We then test the generality of our findings in two realistic spin glass models representing NP-hard optimization problems, the Max Cut and the $3$-XORSAT. First, we demonstrate the same limited efficiency of the first- and second-order local expansion via the ground state gap and the final fidelity. Next, we introduce a yet further enhanced version of the QBCD Hamiltonian by sparsifying it to the density of the local expansion. Remarkably, this exponentially reduced non-locality still suffices to achieve a comparable adiabatic speedup as the full QBCD. These features significantly ease the challenges of implementation, making the approach promising for both experimental platforms and large-scale classical simulations. Finally, we test the COLD approach in the Max Cut problem and find that, even when implemented along optimized annealing paths, its advantage over local expansion quickly diminishes with increasing $L$, whereas the sparsified QBCD consistently outperforms it. In addition, the numerical effort required by COLD grows rapidly, restricting its applicability to very small system sizes ($L\lesssim 10$), at variance with our approach.
@@ II Minimal model of spin-glass bottleneck
From the pictorial representation in Fig. 1a, it can be seen that Eq. (1) describes an Ising chain with uniform ferromagnetic couplings, except for three weaker couplings placed in diametrically opposite positions: two weak ferromagnetic ones ($J$) next to each other, and one weaker antiferromagnetic one ($J^{\prime}$) on the other side. The condition $J^{2}<J^{\prime}$ in Eq. (3) is more technical, and is explained in App. B.4.
The Hamiltonian in Eq. (1) posseses both $\mathbb{Z}_{2}$ fermion parity and $\mathbb{Z}_{2}$ reflection parity symmetry:
$\Pi_{\mathrm{F}}=\prod_{j=1}^{L}\sigma_{j}^{x},\quad\big[\Pi_{\mathrm{F}},H[\lambda]\big]=0;$ (4)
$\Pi_{\mathrm{R}}\sigma_{j}^{\alpha}\Pi_{\mathrm{R}}=\sigma_{L-j+1}^{\alpha},\quad\left[\Pi_{\mathrm{R}},H[\lambda]\right]=0,\quad\alpha=x,y,z,$ (5)
The second one is highlighted with a grey dashed line in Fig. 1a. In the following, we focus on the even fermion parity sector, dynamically accessible from the ground state at $\lambda=0$.
@@ II.1 Simple limits
[…]
The ground state undergoes an exponentially small avoided crossing with the first excited state at an intermediate value of $\lambda$, denoted by $\lambda_{c}$. Its explicit expression is discussed in App. B.4. Across this critical value, the ground state changes from a dressed version of the states in Fig. 1c–d ($\lambda<\lambda_{c}$) to a dressed version of the state in Fig. 1b ($\lambda>\lambda_{c}$), involving a macroscopic flipping of half the chain. This behavior resembles transitions found in spin-glass models.
In the remainder of this section, we provide a thorough description of the avoided crossing, correcting minor omissions in Ref. [99] while preserving the key results.
@@ II.2 Fermionic representation and reflection parity symmetry
By means of a Jordan-Wigner transformation, the spin chain Eq. (1) can be mapped to non-interacting free fermions, as shown in App. A.1). Upon introducing a Majorana fermion representation and by a further transformation using fermionic operators accounting for reflection symmetry (see App. A.2 and App. A.4, respectively), one arrives at a one-dimensional hopping model with on-site potentials at the ends of the chain,
@@
where the $\Gamma$ operators satisfy fermionic anti-commutation relations and possess reflection parity symmetry:
$\{\Gamma_{k}^{+},\Gamma_{k^{\prime}}^{-}\}=\delta_{k,k^{\prime}},\qquad\{\Gamma_{k}^{+},\Gamma_{k^{\prime}}^{+}\}=\{\Gamma_{k}^{-},\Gamma_{k^{\prime}}^{-}\}=0,$ (7a)
$\widetilde{\Pi}_{\mathrm{R}}\Gamma_{k}^{\pm}\widetilde{\Pi}_{\mathrm{R}}=\pm i\Gamma_{k}^{\mp},\qquad(\Gamma^{\pm}_{k})^{\dagger}=\Gamma^{\mp}_{k}.$ (7b)
These identities make use of the chiral reflection parity operator $\widetilde{\Pi}_{\mathrm{R}}$, which is discussed in detail in App. A.3.
Introducing the vectors
@@
Thus, diagonalizing $H[\lambda]$ reduces to diagonalizing the simpler one-particle hopping matrix $\mathcal{H}[\lambda]$. Here, calligraphic letters denote operators on this effective space, and we use $|\psi)$ for the corresponding single-particle modes.
@@ II.3 Description of the avoided crossing
Although the matrix $\mathcal{H}[\lambda]$ in Eq. (9) cannot be diagonalized generally, the entire spectrum can be studied by a Fourier ansatz, due to its translational invariance in the bulk. Thus, the eigenstates must have a plane (or evanescent) wave structure in the bulk, and the on-site potentials acting on the ends of the chains only affect the boundary conditions, as in the case of a piecewise constant (or delta-function) potential.
Figure 2: (a) Sketch of the boundary-localized eigenmodes $|\psi_{\mathrm{L}})$$|\psi_{\mathrm{R}})$ of the effective hopping Hamiltonian (9). (b) Spectrum of the hopping Hamiltonian (9) for $\ell=8$, $J=0.5$, $J^{\prime}=0.27$. The many-body ground state of the $\Gamma$ Hamiltonian in Eq. (6) is formed by populating all the negative-energy modes at $\lambda=0$ (solid lines) and leaving the positive-energy modes empty (dashed lines). The two localized modes $\epsilon_{\mathrm{L}}$ (blue) and $\epsilon_{\mathrm{R}}$ (orange) undergo an avoided crossing at $\lambda=\lambda_{\mathrm{c}}$: while $\epsilon_{\mathrm{R}}$ contributes to the ground state for $\lambda<\lambda_{\mathrm{c}}$ (continuous portion of the orange line), it is replaced by $\epsilon_{\mathrm{L}}$ for $\lambda>\lambda_{\mathrm{c}}$ (the continuous portion of the blue line). Since the spatial overlap between $|\psi_{\mathrm{L}})$ and $|\psi_{\mathrm{R}})$ is exponentially small in $L$, also the gap $\Delta_{\mathrm{min}}$ remains exponentially small according to Eq. (12).
As shown in App. B, the spectrum of the effective hopping model, Eq. (9), is essentially composed of plane wave eigenmodes. Under the condition $J^{2}<J^{\prime}$, however, a couple of boundary modes are present (see also Eq. (94)): $|\psi_{\mathrm{L}})$, localized at $j=0$, and $|\psi_{\mathrm{R}})$, at $j=L$. They are characterized by the exponential decay,
@@
see Fig. 2a. Their eigenvalues undergo an avoided crossing at $\lambda=\lambda_{c}$ with an exponentially small gap
$\Delta_{\mathrm{min}}\propto e^{-\alpha L},\qquad\alpha=\frac{1}{2}\ln\frac{J^{\prime}(1-J^{2})}{J^{2}-(J^{\prime})^{2}}>0,$ (12)
where $\lambda_{c}$ depends on $J$ and $J^{\prime}$; see App. B.4. This gap limits adiabatic driving, since the ground state switches from $|\psi_{\mathrm{R}})$ to $|\psi_{\mathrm{L}})$ across $\lambda_{c}$ (Fig. 2b), requiring the transfer of a particle across the chain, i.e., flipping an extensive number of spins.